bill is 15 years younger than eric. together they are 37 years old. how old is bill and eric?
Answer by macston(5194)

.
B=Bill's age; E=Eric's age=B+15yrs
.
B+E=37years
B+(B+15yrs)=37yrs
2B=22yrs
B=11yrs
ANSWER 1: Bill is 11 years old.
E=B+15yrs=11yrs+15yrs=26yrs
ANSWER 2: Eric is 26 years old.